### v1

@@ XI-C Numerical Evidence showing LLMAs are RIBUM
We perform the max-margin based utility reconstruction from Algorithm 2 for the product quality identification task and illustrate the reconstructed utilities in Figure 13.
For illustration purposes, we consider recommend, durability and expectations as the actions, hence making the action space $|\mathcal{U}|=8$. To show our methods can be used with black-box LLMs, we use ChatGPT-4o-mini for this experiment. The state space is $\mathcal{X}=\{0=\textit{bad \ product},1=\textit{good \ product}\}$. We consider the different environments to have different ratings; i.e., the people who rate products at 11 11 are considered as part of a single environment. The utilities can be used to interpret the behavior of the LLM and can be used in lieu of explicit utilities of the form (33).
@@ XI-D Herding In LLMAs
Figure 14: Public prior freezes in finite time when LLMAs learn using social learning protocol of Algorithm 4 for detecting hate-speech peddlers. The time taken (average over 1010 1010 runs) to form an information cascade is different for different values of the initial public prior $\pi_{0}(x=0)$ and different true underlying states. Note that for all initial priors, only a few observations are enough if the underlying state is 00 0.
@@
We first study the freezing of public prior when an information cascade (Definition 1) happens in a sequence of LLMAs performing Bayesian inference for hate-speech classification. We consider 8080 8080 timepoints and the results average over 1010 1010 runs. The result is illustrated in Figure 14 for different values of initial public prior. In each subplot, the different lines correspond to the different true underlying states. It can be seen that the public prior for state 0 converges to higher values as the initial prior for state 0 is increased. Hence, the initial prior decides what the prior will freeze.
We show, for a sample path realization, how LLMAs herd in the actions that they take in Figure 15. Each of the subplots corresponds to a different initial public prior, and the lines within each subplot are for a different underlying true state. The reason why it is more resilient to herding when in state 4 is that there is a stronger observation likelihood. All the other states, for a strong enough prior (0.520.52 0.520.52), herd to an incorrect state (00 0). Hence classifying an HSP as a non-HSP.
Figure 16 illustrates the regions of herding a case with 33 33 states. The true underlying state in this simulation is 00 0, and therefore, it can be seen that most of the initial public prior corresponds to predicting state 00 0. We assume an observation matrix with 33 33 observations, with an observation matrix $B=[0.7,0.2,0.1;0.1,0.7,0.2;0.2,0.1,0.7]$ and assume an identity utility function for the LLMAs. However, the top-left and bottom-left corners of the triangle show that herding to the wrong state can happen if the public prior is too strong.
Figure 16: Regions of herding in LLMAs for Bayesian social learning with $|\mathcal{X}|=3$ states. Even though the underlying state is $x=0$, in the bottom-left and top-left regions, the actions are 22 22 and 11 11, respectively, because of a strong prior.
@@ XI-E Optimal Stopping for delaying herding in LLMAs
Figure 18: Incentivized stochastic control of autonomous LLMAs. The public prior (left) converges slower as a result of incentivization. Also note that the incentive (right) is a supermartingale (as proved in Theorem 5). This is for a policy of the form (35) threshold parameter of $\theta=0.4$.
@@
In Figure 17, we perform a parameter sweep over $\theta\in[0,1]$ with a grid cardinality of 100100 100100, and evaluate three different aspects. The results are averaged over 100100 100100 independent runs. In the top-left corner, we evaluate the cost function as a function of the threshold parameter. We observe that the cost is minimum for a higher parameter value. And a higher parameter corresponds to a smaller stop time. The accuracy is computed as the proportion of times the correct state is predicted. The sudden jumps in the cost are due to the fact that a small change in the policy parameter is enough to shift the public prior shifts from a region of learning to herding.
@@
We first illustrate our result of Theorem 5 in Figure 18, where it can be seen that the sequence of incentives is a submartingale sequence. We fix the threshold parameter of (35) to $\theta=0.4$. The public prior converges much slower due to the modified cost function (the additional incentive term in 23); hence, using our stopping time formulation, we can extract more private observations, which can be later used to get more accurate estimates.
Figure 19: The total cumulative incentive spent (bottom) and the achieved classification rate (top) for different values of threshold parameter $\theta$. The policy threshold of around 0.70.7 0.70.7 achieves a tradeoff between the classification rate and the total incentive expenditure. The incentives are constant after a particular parameter because of herding.
Figure 19 shows the value of the cost for the different values of the policy threshold $\theta$. We maintain the same system parameters as the previous two experiments. It can be seen that both the classification rate and the total incentive go up as the policy threshold increases in this example. The sudden jump in the incentive is again due to the sudden switch between the prior region from learning to herding. However, also note that the threshold around 0.6 has an incentive in the range [5,10], and the classification rate is still decent (0.9-1), therefore showing how there is an optimal threshold that optimally achieves a tradeoff between the total incentive expenditure and classification performance.
XI-F Convergence of Stochastic Approximation
Finally, we show how the proposed stochastic approximation algorithm (summarized in Algorithm 7) can be used to optimally estimate the policy parameters for optimal stopping in incentivized LLMAs with system parameters from the previous section. We run 100100 100100 iterations, each of which uses a cost function averaged over 100100 100100 independent runs. We set a linearly decreasing step size from $\beta=0.05$ to 0.0050.005 0.0050.005 and set the parameter perturbation as $\delta=1$ and the approximation factor for the sigmoidal policy of (29) as $\varepsilon=0.3$. The discount factor is $\rho=0.99$. Figure 20
Figure 20: The approximate total incentives corresponding to parameters of a stochastic approximation for estimating the optimal threshold parameters of (26). The incentive of the final iterate is much smaller than the incentive expenditure of a too-small or too-large parameter from Figure 19.
Figure 20 presents how the total incentive expenditure changes with each policy parameter update. It can be seen that after a few iterations, the iterates converge to parameter value such that the incentive expenditure is in the range [5,15], which is close to the optimal incentive from Figure 19.
@@ XII Discussion and Future Work
With the rapid adoption of LLMs across science and engineering, and the emergence of LLM-based agents for automating diverse workflows[76], it is crucial to systematically study their behavior and interactions. This work addresses this need by proposing interpretable models and stochastic control algorithms to engineer systems of LLMAs that perform Bayesian inference across a range of applications. Our study highlights the versatility of LLMAs across various fields, aiming to inspire practitioners from diverse disciplines to explore and integrate these tools into their workflows.
@@ XII-A Summary
We first presented a Bayesian sensor model for constructing a large language model agent (LLMA). This model was a composition of a) an LLM, which was used as an interpretable and configurable sensor for high-dimensional data, and b) a Bayesian engine to represent and update the belief of the underlying state. Second, we put forth the necessary and sufficient conditions for a LLMA to be a rationally inattentive Bayesian utility maximizer (RIBUM). We present algorithms for estimating the utility function of a LLMA  which is RIBUM. The reconstructed utility naturally leads to the interpretability of the actions of the LLMA. These methods are applicable to both the Bayesian sensor model of LLMA we propose and off-the-shelf LLMAs and LLMs. Thirdly we look at Bayesian social learning a sequence of LLMAs. We show that this sequence of agents herd in their actions even if they share their private observations. Then, we formulate optimal stopping problems for failure state detection, which optimally delays herding and improves estimation by allowing the private observations to be shared. The optimal stopping problem is formulated both for a case when the LLMAs are centrally controlled and when they are autonomous. We show that under relatively mild conditions on the observation matrix and the cost function of the LLMA  the optimal policy for the optimal stopping problem has a threshold structure. A stochastic approximation is proposed to estimate the optimal policy efficiently without the knowledge of system parameters. Finally, we show numerical experiments demonstrating the various claims and frameworks proposed in the study.
@@ XII-B Insights
    The Bayesian revealed preferences framework is a systematic way to obtain utility (or cost) functions for black-box LLMAs. It’s especially useful when the state and action space are small, but the LLMA operates in tens of different environments.
@@
    Finally, our numerical experiments are done on real-world datasets but are limited to public textual data. Using vision-language models, the framework proposed in this paper can be used for more sophisticated bio-medical data (composed of images, audio, etc.).
@@ XII-C Limitations
The Bayesian sensor model developed for a LLMA requires domain knowledge to effectively design the prompts (features) that are input into the LLM, producing a low-dimensional output; this knowledge may potentially be sourced from a larger LLM. However, the Bayesian revealed preference framework becomes computationally infeasible in scenarios with large state and action spaces. Our formulation of stopping time and the structural results presented are focused solely on detecting a single state, specifically for identifying a failure state. While our findings are derived from the task of Bayesian inference, LLMAs in practice can undertake more complex tasks, indicating a need to extend this work to more general settings.
@@ XII-D Research Directions and Future Work
With technological innovation in LLMs, many players in the market have launched their own LLM agents. Such agents often have more agency than those modeled in this paper, and hence, controlling the behavior of these agents would be more challenging. Further, these agents will inevitably interact with each other through the content they generate or on online platforms. Such interaction would lead to a Bayesian update in their beliefs either explicitly by design as modeled in the paper or implicitly as these agents improve. It is therefore important to consider the models presented in this study in future and related research. Given the rise of LLMAs, we find the following four research directions particularly exciting.
@@ XII-D1 Applicability in other paradigms
XII-D2 Search for a unified interpretable models
The interpretable models presented in this paper function as black-box models for the LLMAs. However, significant research has focused on glass-box models that leverage mechanistic interpretability or utilize explainable features for transformers, the neural architecture underlying LLMs. Integrating these two approaches into a unified model would be advantageous, allowing for explanations of both the LLMA interactions with their external environment and the reasoning behind their behavior. Such a framework could be a valuable tool for analyzing various challenges related to the real-world deployment of LLMAs, including reliability and safety.
XII-D3 Network of LLMAs and Data Fusion
The LLMAs in this study are static, homogeneous, and arranged in a line graph. However, a more generalized network of LLMAs could be deployed as Bayesian sensors for distributed state estimation. These LLMA sensors could then perform data fusion using standard techniques, with careful measures to prevent data incest. Additionally, LLMAs could possess asymmetric private observations and be fine-tuned on individualized datasets, creating a network akin to a distributed mixture of millions of experts [26].
XII-D4 Human In The Loop with LLMAs
In real-world applications, LLMAs frequently interact with humans to receive feedback that can greatly enhance task performance [68]. This necessitates advanced models that account for human-agent interactions to better align cost functions, communication protocols, and sensing mechanisms. As these LLMAs become more widely deployed across various applications, they will form natural networks with humans. Investigating how these LLMAs can share private information both reliably and securely to enhance usability and efficiency presents a compelling and open research challenge.
@@ A-A Proof of Theorem 1
    $\mathds{P}(x|u)=\sum_{y\in\mathcal{Y}}\frac{p_{m}(x,y,u)}{p_{m}(u)}=\sum_{y\in\mathcal{Y}}p_{m}(y|u)p_{m}(x|y)$ (36)
@@
    $B(m)^{\prime}(u|x)=\mathds{P}_{m}(u|x)=\sum_{y\in\mathcal{Y}}p_{m}(u|y)B(m)(y|x)$ (37)
@@
    $\displaystyle Z(\mathds{P}(u|x))$
    $\displaystyle=\max_{m\in\mathcal{M}}~{}c_{m}+J(\mathds{P}(u|x),r_{m})-J(\mathds{P}_{m}(u|x),r_{m})$
    $\displaystyle\implies J(\mathds{P}_{l}(u|x))-c_{l}$
    $\displaystyle\geq J(\mathds{P}(u|x))-C(\mathds{P}(u|x)),~{}\forall~{}\mathds{P}(u|x)$
@@ A-D Proof for Theorem 3
The stopping time problem with the socialistic cost of (19) can be decomposed into two cost terms, each corresponding to the cost terms of the stopping time problem for partially observed Markov decision processes of Theorem 12.3.4 of [40].
$\displaystyle\begin{split}C(\pi,1)&=\frac{1}{1-\rho}\min_{u}c_{u}^{\prime}\pi\\ C(\pi,2)&=\sum_{y\in\mathcal{Y}}c_{y}^{\prime}B_{y}\pi+(d+(1-\rho)\Upsilon)e_{1}\pi-(1-\rho)\Upsilon\end{split}$ (43)
@@
    (F2) $P$ is totally positive of order 2 (TP2)

### v2

@@ XI-C Numerical Evidence showing LLMAs are RIBUM
We perform the max-margin-based utility reconstruction using Algorithm 2 when the agent optimizes a utility in order to achieve product quality identification and illustrate the reconstructed utilities in Figure 13.
For illustration purposes, we consider recommend, durability and expectations as the actions, hence making the action space $|\mathcal{U}|=8$. To show our methods can be used with blackbox LLMs, we use ChatGPT-4o-mini for this experiment. The state space is $\mathcal{X}=\{0=\textit{bad \ product},1=\textit{good \ product}\}$. We consider the different environments to have different ratings; i.e., the people who rate products at 11 11 are considered as part of a single environment. The utilities can be used to interpret the behavior of the LLM and can be used in lieu of explicit utilities of the form (33).
@@ XI-D Herding In LLMAs
Figure 14: Public prior freezes in finite time when LLMAs learn using social learning protocol of Algorithm 4 for detecting hate-speech peddlers. The time taken (average over 1010 1010 runs) to form an information cascade is different for different values of the initial public belief $\pi_{0}(x=0)$ and different true underlying states. Note that for a public prior $\pi_{0}(x=0)\geq 0.41$ only a few observations are enough for the LLMAs to herd to a strong belief on the wrong state (state 00 0) regardless of the true underlying state.
@@
We first study the freezing of public belief when an information cascade (Definition 1) happens in a sequence of LLMAs performing Bayesian inference for hate-speech classification. We consider 8080 8080 timepoints, and the results average over 1010 1010 runs. The result is illustrated in Figure 14 for different values of initial public belief. In each subplot, the different lines correspond to the different true underlying states. It can be seen that the public belief for state 0 converges to higher values as the initial prior for state 0 is increased. Hence, the initial prior decides what the prior will freeze.
We show, for a sample path realization, how LLMAs herd in the actions that they take in Figure 15. Each of the subplots corresponds to a different initial public belief, and the lines within each subplot are for a different underlying true state. The reason why it is more resilient to herding when in state 4 is that there is a stronger observation likelihood. All the other states, for a strong enough prior (0.520.52 0.520.52), herd to an incorrect state (00 0). Hence classifying an HSP as a non-HSP.
Figure 16 illustrates the regions of herding a case with 33 33 states. The true underlying state in this simulation is 00 0, and therefore, it can be seen that most of the initial public belief corresponds to predicting state 00 0. We assume an observation matrix with 33 33 observations, with an observation matrix $B=[0.7,0.2,0.1;0.1,0.7,0.2;0.2,0.1,0.7]$ and assume an identity utility function for the LLMAs. However, the top-left and bottom-left corners of the triangle show that herding to the wrong state can happen if the public belief is too strong.
Figure 16: Regions in belief space where LLMAs herd during Bayesian social learning with $|\mathcal{X}|=3$ states. Even though the underlying state is $x=0$, in the bottom-left and top-left regions, the actions are 22 22 and 11 11, respectively, because of a strong public belief.
@@ XI-E Optimal Stopping for delaying herding in LLMAs
Figure 18: Incentivized stochastic control of autonomous LLMAs. The public belief (left) converges slower as a result of incentivization. Also note that the incentive (right) is a supermartingale (as proved in Theorem 5). This is for a policy of the form (35) threshold parameter of $\theta=0.4$.
@@
In Figure 17, we perform a parameter sweep over $\theta\in[0,1]$ with a grid cardinality of 100100 100100, and evaluate three different aspects. The results are averaged over 100100 100100 independent runs. In the top-left corner, we evaluate the cost function as a function of the threshold parameter. We observe that the cost is minimum for a higher parameter value. And a higher parameter corresponds to a smaller stop time. The accuracy is computed as the proportion of times the correct state is predicted. The sudden jumps in the cost are due to the fact that a small change in the policy parameter is enough to shift the public belief shifts from a region of learning to herding.
@@
We first illustrate our result of Theorem 5 in Figure 18, where it can be seen that the sequence of incentives is a submartingale sequence. We fix the threshold parameter of (35) to $\theta=0.4$. The public belief converges much slower due to the modified cost function (the additional incentive term in 23); hence, using our stopping time formulation, we can extract more private observations, which can be later used to get more accurate estimates.
Figure 19: The cumulative incentive spent (bottom) and the achieved classification rate (top) for different values of threshold parameter $\theta$. The policy threshold of around 0.70.7 0.70.7 achieves a tradeoff between the classification rate and the cumulative incentive expenditure. The incentives are constant after a particular parameter because of herding.
Figure 19 shows the value of the cost for the different values of the policy threshold $\theta$. We maintain the same system parameters as the previous two experiments. It can be seen that both the classification rate and the cumulative incentive go up as the policy threshold increases in this example. The sudden jump in the incentive is again due to the sudden switch between the prior region from learning to herding. However, also note that the threshold around 0.6 has an incentive in the range [5,10], and the classification rate is still decent (0.9-1), therefore showing how there is an optimal threshold that optimally achieves a tradeoff between the cumulative incentive expenditure and classification performance.
XI-F Illustration of Stochastic Gradient Algorithm
Finally, we show how the proposed policy gradient algorithm (summarized in Algorithm 7) can be used to optimally estimate the policy parameters for optimal stopping in incentivized LLMAs with system parameters from the previous section. We run 100100 100100 iterations, each of which uses a cost function averaged over 100100 100100 independent runs. We set a linearly decreasing step size from $\beta=0.05$ to 0.0050.005 0.0050.005 and set the parameter perturbation as $\delta=1$ and the approximation factor for the sigmoidal policy of (29) as $\varepsilon=0.3$. The discount factor is $\rho=0.99$.
Figure 20: The approximate cumulative incentives corresponding to parameters of a policy gradient for estimating the optimal threshold parameters of (26). The incentive of the final iterate is much smaller than the incentive expenditure of a too-small or too-large parameter from Figure 19.
Figure 20 presents how the cumulative incentive expenditure changes with each policy parameter update. It can be seen that after a few iterations, the iterates converge to parameter value such that the incentive expenditure is in the range [5,15], which is close to the optimal incentive from Figure 19.
@@ XII Discussion and Future Work
With the rapid adoption of LLMs across science and engineering and the emergence of LLM-based agents for automating diverse workflows[87], it is crucial to systematically study their behavior and interactions. This work addresses this need by proposing interpretable models and stochastic control algorithms to engineer systems of LLMAs that perform Bayesian inference across a range of applications. Our study highlights the versatility of LLMAs across various fields, aiming to inspire practitioners from diverse disciplines to explore and integrate these tools into their workflows.
@@ XII-A Summary
We first presented a Bayesian sensor model for constructing a large language model agent (LLMA). This model was a composition of a) an LLM, which was used as an interpretable and configurable sensor for high-dimensional data, and b) a Bayesian engine to represent and update the belief of the underlying state. Second, we put forth the necessary and sufficient conditions for a LLMA to be a rationally inattentive Bayesian utility maximizer (RIBUM). We present algorithms for estimating the utility function of a LLMA  which is RIBUM. The reconstructed utility naturally leads to the interpretability of the actions of the LLMA. These methods are applicable to both the Bayesian sensor model of LLMA we propose and off-the-shelf LLMAs and LLMs. Thirdly we look at Bayesian social learning in a sequence of LLMAs. We show that this sequence of agents herd in their actions even if they share their private observations. Then, we formulate optimal stopping problems for failure state detection, which optimally delays herding and improves estimation by allowing the private observations to be shared. The optimal stopping problem is formulated both for a case when the LLMAs are centrally controlled and when they are autonomous. We show that under relatively mild conditions on the observation matrix and the cost function of the LLMA  the optimal policy for the optimal stopping problem has a threshold structure. A policy gradient algorithm is proposed to estimate the optimal policy efficiently without the knowledge of system parameters. Finally, we show numerical experiments demonstrating the various claims and frameworks proposed in the study.
@@ XII-B Insights
    The Bayesian revealed preferences framework is a systematic way to obtain utility (or cost) functions for blackbox LLMAs. It’s especially useful when the state and action space are small, but the LLMA operates in tens of different environments.
@@
    Finally, our numerical experiments are conducted on real-world datasets but are limited to public textual data. Using vision-language models, the framework proposed in this paper can be used for more sophisticated bio-medical data (composed of images, audio, etc.).
@@ XII-C Limitations
The Bayesian sensor model developed for a LLMA requires domain knowledge to design the prompts (features) that are input into the LLM, producing a low-dimensional output; this knowledge can be obtained from a larger LLM. However, the Bayesian revealed preference framework becomes computationally infeasible in scenarios with large state and action spaces. Our formulation of stopping time and the structural results presented are focused solely on detecting a single state, specifically for identifying a failure state. While our framework is Bayesian inference centric, LLMAs in practice can undertake more complex tasks, indicating a need to extend this work to more general settings. In the Bayesian revealed preferences framework considered in this paper, access to the true action posterior is assumed; however, in practice, only empirical estimates may be available. Additionally, the applications discussed in this work—finance, online content moderation, and product analysis—may limit the generalizability of the framework to other domains.
@@ XII-D Research Directions and Future Work
With technological innovation in LLMs, many players in the market have launched their own LLM agents. Such agents often have more functionality than those modeled in this paper, and hence, controlling the behavior of these agents would be more challenging. Further, these agents will inevitably interact with each other through the content they generate or on online platforms. Such interaction would lead to a Bayesian update in their beliefs either explicitly by design as modeled in the paper or implicitly as these agents improve. It is therefore important to consider the models presented in this study in future and related research. Given the rise of LLMAs, we find the following four research directions particularly exciting.
@@ XII-D1 Applicability in other paradigms
XII-D2 Application in Education and Healthcare
The framework proposed in this paper has broader applicability than content moderation, finance, and opinion mining. We briefly discuss the applications of social learning of LLMAs in education and healthcare. In education, there are many potential use cases of LLMs, including customized evaluation, interactive tutorials, and LLM-assisted group discussions  [23]. Therefore, a sequence of LLMAs can be used to analyze different contents, and studying the control of LLMAs is important to reduce the bias, for e.g., while grading students. In healthcare, LLMAs are already used for patient interaction to provide instructions for self-care, and to route the patient to the appropriate healthcare provider. However, multiple LLMAs can be used to analyze a patient’s medical history and provide useful feedback to the doctor. Recently, authors in [93] have further discussed various opportunities for LLMAs in the healthcare domain.
XII-D3 Search for a unified interpretable models
The interpretable models presented in this paper serve as blackbox models for the LLMAs. However, significant research has focused on glassbox (or transparent) models that leverage mechanistic interpretability or utilize explainable features for transformers, the neural architecture underlying LLMs. Integrating these two approaches into a unified model would be advantageous, allowing for explanations of both the LLMA interactions with their external environment and the reasoning behind their behavior. Such a framework could be a valuable tool for analyzing various challenges related to the real-world deployment of LLMAs, including reliability and safety.
XII-D4 Network of LLMAs and Data Fusion
The LLMAs in this study are static, homogeneous, and arranged in a line graph. We have considered a homogenous society of LLM agents for notational convenience and for ease of analysis. Our framework allows for different LLM agents can have different observation matrices, this can model LLM agents who have different LLMs, prompts or pre-training data. Additionally, LLMAs could possess asymmetric private observations and be fine-tuned on individualized datasets, creating a network akin to a distributed mixture of millions of experts [94]. Another direct extension of our framework accommodates LLM agents with different observation and action spaces, provided that each agent is aware of the observation and action spaces of the preceding agent. A generalized network of LLM agents could also be utilized as Bayesian sensors for distributed state estimation. Furthermore, societies of heterogeneous LLM agents communicating with one another warrant further investigation. These LLMA sensors could then perform data fusion using standard techniques, with careful measures to prevent data incest. This paper has looked at stochastic control methods for controlling herding in a sequence of interacting LLMAs. We have formulated two non-standard POMDPs for the centrally controlled and incentivized case. Further to derive structural results, we have looked at the extreme case where there are only two possible actions. These structural results naturally lead to an efficient distributed policy gradient algorithm for searching the optimal policy. It would be interesting to see if newer results from sociology can be used to analyze and control more sophisticated behavior of LLMAs when interacting with each other in heterogeneous societies.
XII-D5 Human In The Loop with LLMAs
In applications such as finance, healthcare, and content moderation, LLMAs frequently interact with humans to receive feedback that can greatly enhance task performance [95]. This necessitates advanced models that account for human-agent interactions to better align cost functions, communication protocols, and sensing mechanisms. As these LLMAs become more widely deployed across various applications, they will form natural networks with humans. Investigating how these LLMAs can share private information both reliably and securely to enhance usability and efficiency presents a compelling and open research challenge.
@@ A-A Proof of Theorem 1
    $\mathds{P}(x|u)=\sum_{y\in\mathcal{Y}}\frac{p_{m}(x,y,u)}{p_{m}(u)}=\sum_{y\in\mathcal{Y}}p_{m}(y|u)p_{m}(x|y).$ (36)
@@
    $B(m)^{\prime}(u|x)=\mathds{P}_{m}(u|x)=\sum_{y\in\mathcal{Y}}p_{m}(u|y)B(m)(y|x).$ (37)
@@
    $\displaystyle Z(\mathds{P}(u|x))=\max_{m\in\mathcal{M}}~{}c_{m}+J(\mathds{P}(u|x),r_{m})-J(\mathds{P}_{m}(u|x),r_{m})$
    $\displaystyle\implies J(\mathds{P}_{l}(u|x))-c_{l}\geq J(\mathds{P}(u|x))-C(\mathds{P}(u|x)),~{}\forall~{}\mathds{P}(u|x)$
@@ A-D Proof for Theorem 3
The stopping time problem with the social welfare cost of (19) can be decomposed into two cost terms, each corresponding to the cost terms of the stopping time problem for partially observed Markov decision processes of Theorem 12.3.4 of [5].
$\displaystyle\begin{split}C(\pi,1)&=\frac{1}{1-\rho}\min_{u}c_{u}^{\prime}\pi,\\ C(\pi,2)&=\sum_{y\in\mathcal{Y}}c_{y}^{\prime}B_{y}\pi+(d+(1-\rho)\Upsilon)e_{1}\pi-(1-\rho)\Upsilon.\end{split}$ (43)
@@
    (F2) $P$ is totally positive of order 2 (TP2).